OpenStudy (anonymous):

Calculate the # of moles present. 521 mg of germanium

sam (.sam.):

Start by finding the g/mol of germanium using the periodic table, http://www.ptable.com/ Looks like its 72.6g/mol. So, you want to find the number of moles, you can use the fraction method. \[521mg =5.21 \times 10 ^{-1} \text{ g of germanium}\] \[5.21 \times 10 ^{-1} \text{ g of germanium} \times \frac{\text{1 mole of germanium}}{\text{72.6 g of germanium}}\] \[5.21 \times 10 ^{-1} \cancel{\text{g of germanium}} \times \frac{\text{1 mole of germanium}}{72.6 \cancel{\text{ g of germanium}}}=7.17 \times 10^{-3} \text{ moles of germanium}\]